# Temporal Scaling Law for Large Language Models

### v3

@@ 3 Use Case #1 of Temporal Scaling Law: Hyperparameter Selection
Specifically, to select the best data mixture proportion for training a larger 468M/1.2B model, in the retrieval stage, following the prior art Xie et al. (2023), we use a small-scale model to locate a group of data mixture proportion candidates. In the rerank stage, we train all located data proportions on the target LLM for acceptable training tokens, and locate the best one by predicting their final losses with our temporal scaling law.
Experiment Setup. In the retrieval stage, we conduct a grid search on a 58M model and locate the Top-5 candidates. In the rerank stage, we train each candidate for 10B tokens before making the test loss prediction with Eq.11-14 on validation datasets. To validate the data mixture choices by temporal scaling law on the target 468M/1.2B models, we (1) primarily use the perplexity metric on the ID-Val and the OOD-Val, and also (2) follow Touvron et al. (2023a); Rae et al. (2021) and employ 7 popular common sense reasoning benchmarks, including BoolQ Clark et al. (2019), HellaSwag Zellers et al. (2019), OpenBookQA Mihaylov et al. (2018), PIQA Bisk et al. (2020), SIQA Sap et al. (2019), StoryCloze Mostafazadeh et al. (2016), and Winogrande Sakaguchi et al. (2021) for evaluating the final fully trained models. See the appendix H for more experimental details.
Experiment Result. We compare the data mixture proportion selected by our retrieval-rerank pipeline with the following baselines: (1) the Top-1 selected directly on the much smaller model and (2) the Top-1 selected by comparing the real test losses on ID-Val and OOD-Val validation sets after training 10B tokens on 468M/1.2B model. We find that both (1) and (2) make the same choice from the Top-5 candidates, and our pipeline locates a different choice from them. As shown in Tab. 4, our choice achieves lower perplexity across all settings after completing pretraining (i.e., 400B tokens). Additionally, as shown in Tab. 3, our choice yields superior benchmark performance in 12 of 14 metrics. Consequently, our approach of directly selecting hyperparameters on the target LLM is more effective and incurs an acceptable cost, which does not necessitates fully training the model.
@@ 4 Use Case #2 of Temporal Scaling Law: Revisit the Training Strategy
The temporal scaling law has afforded the possibility to study LLM pretraining in a finer granularity. As noted in Sec. 2.3, tokens on different positions possess a fundamental bias in learning difficulty. Specifically, head tokens are generally harder to predict than tail tokens. However, as suggested by Eqs. 6 and 8, $a_{0}$ and $a_{1}$ remain constant after the separation point. Therefore, for $N>N_{sep}$:
@@
which is unrelated to token position $i$, suggesting that LLMs learn equally on different token positions. To validate this observation, we trace the actual loss decrease observed during training along different token positions. It is observed that in a later training stage where the separation point conditions in Eq. 7 are already fulfilled, test loss decrease among all token positions tends to be uniform across all settings (see Fig. 11 for results). Such a consistency between theoretical prediction and experimental observation strongly validates the reasonableness and effectiveness of our temporal scaling law. Experiments also show that simply averaging losses on all token positions is an effective strategy for training LLMs. Refer to appendix I for more experimental settings and results.
@@ Appendix B Further Experiment Details
17:  /****** Predict the test loss curve of $M$*****/
@@
Hyperparameters of Models Used. We report the details of model hyperparameters and training hyperparameters in Tab. 5. Note that we use the 9.8M and the 58M models for illustrating our temporal scaling law. Meanwhile, we apply the predictions and further applications of our temporal scaling law to the larger 468M and the 1.2B models.
Parameters for Packages. We report the version numbers of used packages in Tab. 7.
@@ Appendix C Theoretical Insights for the Dynamic Hyperbolic Law
Following prior work on validating scaling laws Kaplan et al. (2020); Hoffmann et al. (2022), we conducted comprehensive experiments and empirically confirmed the effectiveness of the hyperbolic pattern in the main article. In the main article, the main functional forms that we have considered are the logarithmic function ($\mathcal{L}_{i}=a_{1}\log(a_{2}i+a_{3})+a_{4}$), power-law ($\mathcal{L}_{i}=a_{1}i^{a_{2}}+a_{3}$), log-log function ($\mathcal{L}_{i}=a_{1}\log(a_{2}\log(i)+a_{3})+a_{4}$), and exponential function ($\mathcal{L}_{i}=a_{1}\cdot a_{2}^{i}+a_{3}$) (see Sec. 2.4 for the experimental comparisons with those functions). In addition, we provide a theoretical justification for selecting the hyperbolic function among candidate forms. Specifically, the loss with respect to token positions should approach a lower bound, since the cross-entropy loss has a minimum of zero. This constraint limits the valid choices to the power-law, exponential, and hyperbolic functions. We then analyze the simplified forms of the power-law and exponential functions: for the power-law, $a_{1}>0$, $a_{2}<0$, $a_{3}=0$; for the exponential, $a_{1}>0$, $10<a_{2}<$ 1, $a_{3}=0$. It is straightforward to show that the exponential function decays faster than the power-law. Formally, for all valid parameters, there exists an $i_{0}\in\mathbb{N}$ such that $\mathcal{L}^{power}_{i}>\mathcal{L}^{exp}_{i}$ for all $i>i_{0}$. Given that real-world corpus data typically does not yield losses approaching zero, the exponential function is impractical. Lastly, we note that the hyperbolic function can be rewritten in a power-law-like form as $\mathcal{L}_{i}=a_{0}(a_{1}i+1)^{-1}+a_{2}$. Since this formulation not only satisfies theoretical constraints but also yields better empirical performance, we adopt it as the underlying structure for modeling token position loss.
@@
Appendix D Pipeline for Predicting Training Outcomes Using the Temporal Scaling Law
@@
We provide the pipeline for predicting the training outcomes using the temporal scaling law in algorithm 1. Note that we use non-linear least squares to solve for all fitting parameters. Normally, the fitting would converge in $<10^{3}$ steps, which takes <1 second on a CPU. We conduct a re-fit for $a_{0}$ and $a_{1}$ to mitigate the effect of fluctuations (as observed in Fig. 2 in the article).
Appendix E Complete Fitting Results
@@
Due to the page limit, we omitted some fitting results when illustrating our proposed temporal scaling law. In this section, we will present those fitting figures in detail.
E.1 More Results of Dynamic Hyperbolic-Law
@@
As a supplement to Figure 1 in Section 3.3 of the article, we provide fitting results of the dynamic hyperbolic-law for both the 9.8M and 58M models after training for 100B, 200B, 300B, and 400B tokens and on the OOD-Val in Fig. 6 and Fig. 7, respectively. Those results indicate that our dynamic hyperbolic-law consistently achieves accurate fitting results across model sizes and training steps.
E.2 More Results of Temporal Scaling Law
@@
As a supplement to Figure 2 in Section 3.3 of the article, we provide the results on the OOD-Val in Fig. 8. We also provide fitting results on both the ID-Val and the OOD-Val for the larger 468M, 1.2B, and 6.7B model scales in Fig. 9 and Fig. 10. Across different settings on model scales and validation set distributions, our temporal scaling law is able to depict the general trend of parameter evolution and reduce the influence of fluctuations as much as possible.
Model LAMBADA Wikitext
9.8M 0.9887 0.9809
58M 0.9805 0.9899
468M 0.9831 0.9920
1.2B 0.9859 0.9930
6.7B 0.9872 0.9927
@@
Figure 4: More temporal scaling law fitting results for the 468M GPT-NeoX model and the 1.4B MoE model on both the ID-Val and OOD-Val. Note that the fitting for $a_{1}$ on the MoE model has smaller $R^{2}$ values than the fit for other parameters. This is due to the outliers in the beginning of the graph, as we applied a smaller number of total training tokens for the model, and the training was not stablized in the beginning. Nevertheless, with the outlier filting mechanism described in Sec. 2.3 and algorithm 1, the temporal scaling law successfully portraits the overall trend of $a_{1}$.
@@
E.3 Fitting Results on More Datasets
In the article, we have proposed the temporal scaling law to precisely fit the temporal evolution of test loss during LLM pretraining. To better demonstrate the validity of our temporal scaling law, we also validate the fitting results of the 9.8M and 58M models on the LAMBADA Paperno et al. (2016) and the WikiText Merity et al. (2017) datasets. Specifically, we use the validation and the test splits for the LAMBADA dataset, and use the wikitext-2-v1 domain for the WikiText dataset. The results are shown in Tab. 8. We also provide the fitting results for the larger 468M, 1.2B, and 6.7B models. As shown in Tab. 8, all results achieve $R^{2}>0.98$, well demonstrating the effectiveness of our temporal scaling law.
@@ E.3 Fitting Results on More Datasets
Appendix F Generalization to More Model Structures
@@
Experimental Setup. The 468M GPTNeoXForCausalLM model shares the same hyperparameters with the 468M LLaMA-based model listed in Tab. 5. The 1.4B MoE model is built based on the LLaMA architecture with 16 total experts and adopts a Top-2 activation strategy, following the common setting in various MoE research Su et al. (2024a, b); Huang et al. (2024). Both models are pre-trained on the Pile dataset used in the main article. Specifically, the 468M GPTNeoXForCausalLM model is pre-trained with 200B tokens, and the 1.4B MoE model is pre-trained with 100B tokens. Detailed settings are listed in Tab. 6.
Experiment Results. We present the fitting results on both ID-Val and OOD-Val of the two pre-trained models in Fig. 4. The proposed temporal scaling law successfully genralizes to those architectures, capturing the primary patterns of parameter evolution for the GPT-NeoX and the MoE models, well demonstrating its generalizability.
Appendix G Test Loss Prediction Results on OOD-Val
In Fig. 3, we presented the test loss prediction results on the ID-Val for the scaled-up larger models. Additionally, we present the prediction results on the OOD-Val in Fig. 5. Same with results on the ID-Val, our temporal scaling law consistently generates reliable results over the comparing baselines.
Appendix H More Details for Use Case #1: Hyperparameter Selection
Due to the page limit, we omitted some details for the use case #1 in Section 4 in the article. They are thoroughly described below for reproducibility.
@@ Appendix H More Details for Use Case #1: Hyperparameter Selection
Proportion Candidate Generation. In the retrieval stage, we generate a group of proportion candidates by conducting grid search on domain weights in the Pile dataset Gao et al. (2021). Specifically, we first calculate the original domain weight according to the original epoch settings noted by Gao et al. (2021) on our applied 32k tokenizer. Note that this is also the domain weight applied in Section 3.2 of the article. For the grid search, we adopt a simple but popular pipeline by selecting a domain, modifying its domain weight by a random value $s\in[-0.05,0.05]$, and normalizing the domain weight in other domains.
@@
Appendix I More Details for Use Case #2: Revisit the Training Strategy
@@
In Section 5 of the article, we have stressed that a fundamental bias in learning difficulty based on token position exists. Specifically, head tokens (with shorter context) are generally harder to predict than tail tokens (with longer context), due to the increased uncertainty caused by limited contextual information available for earlier tokens in a sequence. Surprisingly, our temporal scaling law suggests that LLMs learn equally on different token positions after an early training period (as shown by Equation 17 in the article), despite the learning difficulty bias.
@@
Observation for Actual Loss Decrease on Positions. We use the larger 468M and the 1.2B models to validate our suggestion. To validate this observation, we plot the actual loss decrease pattern observed during training along different token positions in Fig. 11. As shown in Fig. 11, in the early training period of “from 20B to 40B tokens”, the head tokens suffer from less loss decrease due to a higher learning difficulty. However, after the early training period, in a later period of ”from 140B to 160B tokens” that the separation point conditions in Equation 7 in the article are already fulfilled, test loss decrease among all token positions tends to be uniform across all settings. Therefore, we can infer from this observation that the learning dynamics derived from our temporal scaling law are authentically presented in LLM pre-training.
@@
Insights for Weighting Strategies. We hypothesize that the default strategy for training generative language models, in which losses on tokens in all positions are simply averaged, is an effective solution. To validate the hypothesis, we conduct LLM pretraining on the larger 468M and the 1.2B models with three position-based weighting strategies: Head suppression, Body suppression, and Tail suppression, applied only after the $N_{sep}$ point. The implementation details of these strategies are described in Tab. 9. Note that for the Suppression strategies, we normalize the weighted losses to ensure the average weight for each token position is 1.0, and thus make the corresponding average loss comparable with the default practice. We apply the pre-training settings as in Section 3.2 of the article for comparing different position-based weighting strategies. As shown in Tab. 10, despite being weighted on different token positions, weighting different tokens by position during LLM pretraining probably cannot yield better results than the default practice, though different positions possess fundamental bias in learning difficulty.
@@
To further validate that the default practice actually trains a competitive model compared to the weighting strategies, we test the model performance on the Common Sense Reasoning benchmark described in Section 4 of the article, and report average model performance on 0-shot, 1-shot, and 5-shot settings. As shown in Tab. 11, all position-based weighting strategies acquire comparable or slightly inferior average results to the default practice, in which no weighting strategies are attached. On individual tasks, the default practice even achieves top-2 accuracies among 11 of 14 settings, surpassing all weighting variants. This indicates that weighting different tokens by position during LLM pre-training probably cannot yield better results than the default training strategy, further demonstrating that it is unnecessary to re-weight tokens by their positions during LLM pre-training.
@@
Figure 11: Test loss decrease on different token positions in the given training period for the 468M and the 1.2B models. “$\Delta_{20B-40B}$” means the decrease of loss at each token position from “being trained with 20B tokens” to “being trained with 40B tokens” etc. After an early training period, the loss decrease tends to be uniform across all token positions (i.e., $\Delta_{140B-160B}$).

### v4

@@ 3 Use Case #1 of Temporal Scaling Law: Hyperparameter Selection
Specifically, to select the best data mixture proportion for training a larger 468M/1.2B model, in the retrieval stage, following the prior art Xie et al. (2023); Feng et al. (2024), we use a small-scale model to locate a group of data mixture proportion candidates. In the rerank stage, we train all located data proportions on the target LLM for acceptable training tokens, and locate the best one by predicting their final losses with our temporal scaling law.
Experiment Setup. In the retrieval stage, we conduct a grid search on a 58M model and locate the Top-5 candidates. In the rerank stage, we train each candidate for 10B tokens before making the test loss prediction with Eq.11-14 on validation datasets. To validate the data mixture choices by temporal scaling law on the target 468M/1.2B models, we (1) primarily use the perplexity metric on the ID-Val and the OOD-Val, and also (2) follow Touvron et al. (2023a); Rae et al. (2021) and employ 7 popular common sense reasoning benchmarks, including BoolQ Clark et al. (2019), HellaSwag Zellers et al. (2019), OpenBookQA Mihaylov et al. (2018), PIQA Bisk et al. (2020), SIQA Sap et al. (2019), StoryCloze Mostafazadeh et al. (2016), and Winogrande Sakaguchi et al. (2021) for evaluating the final fully trained models. See the appendix˜J for more experimental details.
Experiment Result. We compare the data mixture proportion selected by our retrieval-rerank pipeline with the following baselines: (1) the Top-1 selected directly on the much smaller model and (2) the Top-1 selected by comparing the real test losses on ID-Val and OOD-Val validation sets after training 10B tokens on 468M/1.2B model. We find that both (1) and (2) make the same choice from the Top-5 candidates, and our pipeline locates a different choice from them. As shown in Tab.˜5, our choice achieves lower perplexity across all settings after completing pretraining (i.e., 400B tokens). Additionally, as shown in Tab.˜4, our choice yields superior benchmark performance in 12 of 14 metrics. Consequently, our approach of directly selecting hyperparameters on the target LLM is more effective and incurs an acceptable cost, which does not necessitates fully training the model.
@@ 4 Use Case #2 of Temporal Scaling Law: Revisit the Training Strategy
The temporal scaling law has afforded the possibility to study LLM pretraining in a finer granularity. As noted in Sec.˜2.3, tokens on different positions possess a fundamental bias in learning difficulty. Specifically, head tokens are generally harder to predict than tail tokens. However, as suggested by Eqs.˜6 and 8, $a_{0}$ and $a_{1}$ remain constant after the separation point. Therefore, for $N>N_{sep}$:
@@
which is unrelated to token position $i$, suggesting that LLMs learn equally on different token positions. To validate this observation, we trace the actual loss decrease observed during training along different token positions. It is observed that in a later training stage where the separation point conditions in Eq.˜7 are already fulfilled, test loss decrease among all token positions tends to be uniform across all settings (see Fig.˜11 for results). Such a consistency between theoretical prediction and experimental observation strongly validates the reasonableness and effectiveness of our temporal scaling law. Experiments also show that simply averaging losses on all token positions is an effective strategy for training LLMs. Refer to appendix˜K for more experimental settings and results.
@@ 5 Conclusion
Acknowledgment
This work was supported by National Natural Science Foundation of China (Nos. 62525103, 62271281, 62441235) and Beijing Natural Science Foundation (Nos. L247026).
@@ A.2 Scaling Laws for Language Models
Hyperbolic Logarithmic Power-law Log-log Exponential
Eq.˜5 >99% 5% 67% 31% 0%
Eq.˜6 0.690/0.921 0.788/0.659 0.342/0.441 0.953/0.731 -1.231/-5.002
Table 9: Comparison of functional form choices for Eqs.˜5 and 6 on ID-Val for the 58M model. For Eq.˜5, we report the percentage of checkpoints that could achieve $R^{2}>0.95$ for each functional form. For Eq.˜6, we report the $R^{2}$ values fitting with each functional form ($a_{0}$$a_{1}$). Bold represents the best result.
@@ Appendix B Further Experiment Details
17: /****** Predict the test loss curve of $M$ ******/
@@
Hyperparameters of Models Used. We report the details of model hyperparameters and training hyperparameters in Tab.˜6. Note that we use the 9.8M and the 58M models for illustrating our temporal scaling law. Meanwhile, we apply the predictions and further applications of our temporal scaling law to the larger 468M and the 1.2B models.
Parameters for Packages. We report the version numbers of used packages in Tab.˜8.
@@ Appendix C Theoretical Insights for the Dynamic Hyperbolic Law
Following prior work on validating scaling laws Kaplan et al. (2020); Hoffmann et al. (2022), we conducted comprehensive experiments and empirically confirmed the effectiveness of the hyperbolic pattern in the main article. In the main article, the main functional forms that we have considered are the logarithmic function ($\mathcal{L}_{i}=a_{1}\log(a_{2}i+a_{3})+a_{4}$), power-law ($\mathcal{L}_{i}=a_{1}i^{a_{2}}+a_{3}$), log-log function ($\mathcal{L}_{i}=a_{1}\log(a_{2}\log(i)+a_{3})+a_{4}$), and exponential function ($\mathcal{L}_{i}=a_{1}\cdot a_{2}^{i}+a_{3}$) (see Sec.˜2.4 for the experimental comparisons with those functions). In addition, we provide a theoretical justification for selecting the hyperbolic function among candidate forms. Specifically, the loss with respect to token positions should approach a lower bound, since the cross-entropy loss has a minimum of zero. This constraint limits the valid choices to the power-law, exponential, and hyperbolic functions. We then analyze the simplified forms of the power-law and exponential functions: for the power-law, $a_{1}>0$, $a_{2}<0$, $a_{3}=0$; for the exponential, $a_{1}>0$, $0<a_{2}<1$, $a_{3}=0$. It is straightforward to show that the exponential function decays faster than the power-law. Formally, for all valid parameters, there exists an $i_{0}\in\mathbb{N}$ such that $\mathcal{L}^{power}_{i}>\mathcal{L}^{exp}_{i}$ for all $i>i_{0}$. Given that real-world corpus data typically does not yield losses approaching zero, the exponential function is impractical. Lastly, we note that the hyperbolic function can be rewritten in a power-law-like form as $\mathcal{L}_{i}=a_{0}(a_{1}i+1)^{-1}+a_{2}$. Since this formulation not only satisfies theoretical constraints but also yields better empirical performance, we adopt it as the underlying structure for modeling token position loss.
Appendix D Comparison with Other Functional Form Choices
In this article, for fitting general “increasing” and “decreasing” patterns (i.e., Eqs.˜5 and 6), we consider using the basic functional forms described in appendix˜C. We compare the fitting results of different functional forms on the 58M model. Specifically, for Eq.˜5, we report the percentage of checkpoints that could achieve $R^{2}>0.95$ for each functional form. For Eq.˜6, we report the $R^{2}$ values fitting with each functional form. The results are presented in Tab.˜9. Based on these results, we selected the hyperbolic function for fitting the token position loss, the log-log function for $a_{0}$, and the hyperbolic function for $a_{1}$, as these provided the best fit.
@@ Appendix D Comparison with Other Functional Form Choices
Appendix E Pipeline for Predicting Training Outcomes Using the Temporal Scaling Law
@@
We provide the pipeline for predicting the training outcomes using the temporal scaling law in algorithm˜1. Note that we use non-linear least squares to solve for all fitting parameters. Normally, the fitting would converge in $10^{3}$ steps, which takes <1 second on a CPU. We conduct a re-fit for $a_{0}$ and $a_{1}$ to mitigate the effect of fluctuations (as observed in Fig.˜2).
Appendix F Complete Fitting Results
@@
In this section, we present complete fitting figures illustrating our proposed temporal scaling law.
F.1 More Results of Dynamic Hyperbolic-Law
@@
As a supplement to Fig.˜1, we provide fitting results of the dynamic hyperbolic-law for both the 9.8M and 58M models after training for 100B, 200B, 300B, and 400B tokens and on the OOD-Val in Fig.˜6 and Fig.˜7, respectively. Those results indicate that our dynamic hyperbolic-law consistently achieves accurate fitting results across model sizes and training steps.
F.2 More Results of Temporal Scaling Law
@@
As a supplement to Fig.˜2, we provide the results on the OOD-Val in Fig.˜8. We also provide fitting results on both the ID-Val and the OOD-Val for the larger 468M, 1.2B, and 6.7B model scales in Fig.˜9 and Fig.˜10. Across different settings on model scales and validation set distributions, our temporal scaling law is able to depict the general trend of parameter evolution and reduce the influence of fluctuations as much as possible.
@@
Figure 4: More temporal scaling law fitting results for the 468M GPT-NeoX model and the 1.4B MoE model on both the ID-Val and OOD-Val. Note that the fitting for $a_{1}$ on the MoE model has smaller $R^{2}$ values than the fit for other parameters. This is due to the outliers in the beginning of the graph, as we applied a smaller number of total training tokens for the model, and the training was not stablized in the beginning. Nevertheless, with the outlier filting mechanism described in Sec.˜2.3 and algorithm˜1, the temporal scaling law successfully portraits the overall trend of $a_{1}$.
@@
Appendix G Generalization to More Model Structures
@@
Experimental Setup. The 468M GPTNeoXForCausalLM model shares the same hyperparameters with the 468M LLaMA-based model listed in Tab.˜6. The 1.4B MoE model is built based on the LLaMA architecture with 16 total experts and adopts a Top-2 activation strategy, following the common setting in various MoE research Su et al. (2024a, b); Huang et al. (2024). Both models are pre-trained on the Pile dataset used in the main article. Specifically, the 468M GPTNeoXForCausalLM model is pre-trained with 200B tokens, and the 1.4B MoE model is pre-trained with 100B tokens. Detailed settings are listed in Tab.˜7.
Experiment Results. We present the fitting results on both ID-Val and OOD-Val of the two pre-trained models in Fig.˜4. The proposed temporal scaling law successfully genralizes to those architectures, capturing the primary patterns of parameter evolution for the GPT-NeoX and the MoE models, well demonstrating its generalizability.
Appendix H Generalization to More Learning Schedules
In this article, we mainly focus on the cosine decay scheduler, which is adopted by most mainstream LLM structures Touvron et al. (2023a, b); Bai et al. (2023). To provide insights for other schedulers, we have managed to conduct a preliminary experiment with the 468M model using a simple linear learning rate scheduler. Our results indicate that for this setup, replacing the fitting function for the $a_{2}$ parameter with $a_{2}^{N}=\gamma_{4}\cdot N+\gamma_{5}\ (N>N_{sep})$ yields a high overall fit with $R^{2}=0.9947$. This demonstrates a strong correlation between the temporal evolution of $a_{2}$ and the choice of learning rate scheduler, which is also observed in recent works Luo et al. (2025).
Appendix I Test Loss Prediction Results on OOD-Val
In Fig.˜3, we presented the test loss prediction results on the ID-Val for the scaled-up larger models. Additionally, we present the prediction results on the OOD-Val in Fig.˜5. Same with results on the ID-Val, our temporal scaling law consistently generates reliable results over the comparing baselines.
Appendix J More Details for Use Case #1: Hyperparameter Selection
Due to the page limit, we omitted some details for the use case #1 in Sec.˜3. They are thoroughly described below for reproducibility.
@@ Appendix I Test Loss Prediction Results on OOD-Val
Proportion Candidate Generation. In the retrieval stage, we generate a group of proportion candidates by conducting grid search on domain weights in the Pile dataset Gao et al. (2021). Specifically, we first calculate the original domain weight according to the original epoch settings noted by Gao et al. (2021) on our applied 32k tokenizer. Note that this is also the domain weight applied in Sec.˜2.2. For the grid search, we adopt a simple but popular pipeline by selecting a domain, modifying its domain weight by a random value $s\in[-0.05,0.05]$, and normalizing the domain weight in other domains.
@@
Appendix K More Details for Use Case #2: Revisit the Training Strategy
@@
In Sec.˜4, we have stressed that a fundamental bias in learning difficulty based on token position exists. Specifically, head tokens (with shorter context) are generally harder to predict than tail tokens (with longer context), due to the increased uncertainty caused by limited contextual information available for earlier tokens in a sequence. Surprisingly, our temporal scaling law suggests that LLMs learn equally on different token positions after an early training period (as shown by Eq.˜17), despite the learning difficulty bias.
@@
Observation for Actual Loss Decrease on Positions. We use the larger 468M and the 1.2B models to validate our suggestion. To validate this observation, we plot the actual loss decrease pattern observed during training along different token positions in Fig.˜11. As shown in Fig.˜11, in the early training period of “from 20B to 40B tokens”, the head tokens suffer from less loss decrease due to a higher learning difficulty. However, after the early training period, in a later period of ”from 140B to 160B tokens” that the separation point conditions in Eq.˜7 are already fulfilled, test loss decrease among all token positions tends to be uniform across all settings. Therefore, we can infer from this observation that the learning dynamics derived from our temporal scaling law are authentically presented in LLM pre-training.
@@
Insights for Weighting Strategies. We hypothesize that the default strategy for training generative language models, in which losses on tokens in all positions are simply averaged, is an effective solution. To validate the hypothesis, we conduct LLM pretraining on the larger 468M and the 1.2B models with three position-based weighting strategies: Head suppression, Body suppression, and Tail suppression, applied only after the $N_{sep}$ point. The implementation details of these strategies are described in Tab.˜10. Note that for the Suppression strategies, we normalize the weighted losses to ensure the average weight for each token position is 1.0, and thus make the corresponding average loss comparable with the default practice. We apply the pre-training settings as in Sec.˜2.2 for comparing different position-based weighting strategies. As shown in Tab.˜11, despite being weighted on different token positions, weighting different tokens by position during LLM pretraining probably cannot yield better results than the default practice, though different positions possess fundamental bias in learning difficulty.
@@
To further validate that the default practice actually trains a competitive model compared to the weighting strategies, we test the model performance on the common sense reasoning benchmarks described in Sec.˜3, and report average model performance on 0-shot, 1-shot, and 5-shot settings. As shown in Tab.˜12, all position-based weighting strategies acquire comparable or slightly inferior average results to the default practice, in which no weighting strategies are attached. On individual tasks, the default practice even achieves top-2 accuracies among 11 of 14 settings, surpassing all weighting variants. This indicates that weighting different tokens by position during LLM pre-training probably cannot yield better results than the default training strategy, further demonstrating that it is unnecessary to re-weight tokens by their positions during LLM pre-training.
@@
Figure 11: Test loss decrease on different token positions in the given training period for the 468M and the 1.2B models. $\Delta_{20B-40B}$” means the decrease of loss at each token position from “being trained with 20B tokens” to “being trained with 40B tokens” etc. After an early training period, the loss decrease tends to be uniform across all token positions (i.e., $\Delta_{140B-160B}$).